OpenStudy (anonymous):

Locate the foci of the ellipse. Show your work. x^2/36+y^2/11=1 my answer : General ellipse setup Centre (h,k) is known as : x^2 = (x-h)^2, then h = 0 y^2 = (x-k)^2, then k = 0 The centre is (0,0) x^2/36 + y^2/11 = 1 when x= 0 y^2/11 = 1 ; y = 0 When y= 0, x = 0 x^2/36 = 1 ; x = 0 the foci will be along the y=0 axis ... (-c,0) and (c,0) such that b^2 + c^2 = a^2

OpenStudy (anonymous):

@hartnn please check is it incomplete?

OpenStudy (anonymous):

i need to submit this in evening plzz help

hartnn (hartnn):

yeah, u know how will u get the value of c ?

OpenStudy (michele_laino):

please note that: \[a ^{2}=36, b ^{2}=11\] so?

OpenStudy (anonymous):

11 + c^2 = 36 ; solve for c C=5

OpenStudy (michele_laino):

ok!

hartnn (hartnn):

yes, c=5 is correct :)

OpenStudy (anonymous):

what else?

hartnn (hartnn):

focus (-5,0) and (5,0)

hartnn (hartnn):

thats it

OpenStudy (michele_laino):

please, substitute now, your value for c, into the coordinates of focuses

OpenStudy (anonymous):

confused :/

OpenStudy (michele_laino):

you should get as @hartnn well wrote!

OpenStudy (jhannybean):

equation of an ellipse : \(\dfrac{x^2}{a^2} +\dfrac{y^2}{b^2}=1\) To find the foci: \(a^2 -c^2 =b^2\)

OpenStudy (michele_laino):

for example first focus, F_1 is: \[F _{1}=(c,0)=(5,0)\]

OpenStudy (michele_laino):

please, do the same with other focus

OpenStudy (michele_laino):

\[F _{2}=(-c,0)=...\]

hartnn (hartnn):

you had foci as \((\pm c,0)\) you got c=5 so your foci are \((5,0) and (-5,0)\) what else u need ? you're done!